I'm stumped on 2 word problems dealing with probability. Can someone please help me????
Here they are:
1. Suppose a contest offers $1,000 to the person who can guess the winning four digit number. How many possibilities are there?
2. There are 10 women and 8 men in a club. How many different committees of 6 people can be selected from the group if equal numbers of men and women are to be on the committee?
TIA for your help!
Andrew

Answer by stanbon(48568)

1. Suppose a contest offers $1,000 to the person who can guess the winning four digit number. How many possibilities are there?
----------
# of four-digit numbers: 1000-9999= 8999+1 = 9000
-------------------------------------------------------

2. There are 10 women and 8 men in a club. How many different committees of 6 people can be selected from the group if equal numbers of men and women are to be on the committee?
----
6 on the committe requires 3 men and 3 women.
# of groups of 3 women: 10C3 = [10*9*8]/[1*2*3] = 120
# of groups of 3 men : 8C3 = 8*7*6/1*2*3 = 56
------------------
# of possible committees: 120*56 = 6720
============================================
Cheers,
Stan H.